OpenStudy (anonymous):

if f(x) = (5x-7) and g(x) = (x-2), find (fg)(x), then simplify

OpenStudy (zzr0ck3r):

(fg)(x) = f(x)*g(x) = 95x-7)(x-2) = ?

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

do you mean f(gx)?

OpenStudy (zzr0ck3r):

should that say f(g(x))?

OpenStudy (zzr0ck3r):

are you doing composition of functions?

OpenStudy (anonymous):

if so then put g(x) or x-2 in place of all the x's in f(x) giving you 5(x-2) - 7

OpenStudy (anonymous):

he isnt even here anymore lol

OpenStudy (anonymous):

well your answer is 5x-10-7 or 5x-17 enjoy